OpenStudy (anonymous):

What is the a value of the following function? A. 1/3 B. 1/2 C. 3 D. 2

OpenStudy (anonymous):

@Mertsj I got 1/2

OpenStudy (mertsj):

I don't understand the question. What does it mean, the value of the function?

OpenStudy (anonymous):

I have no clue lol

OpenStudy (mertsj):

Their is something wrong. Are you sure you posted the question exactly right? There are many function values shown on that graph.

OpenStudy (anonymous):

yep I posted the question right, i guess it wants us to try and "find" the value

OpenStudy (anonymous):

The "a" value is "2" because the function is symmetric with x=0. Look at x=1, you have 2 for y Look at x=2, you have 8 for y.

OpenStudy (anonymous):

alright then what?:)

OpenStudy (anonymous):

ax^2 + bx + c = y What you really have is y = ax^2 It's the graph of y = 2x^2.

OpenStudy (mertsj):

Ok. tcarroll is right. I completely missed the "a" part in the question. tcarroll010 will help you now. Good luck.

OpenStudy (anonymous):

so im thinking the answer is 3?

OpenStudy (anonymous):

Since the graph is symmetric with x=0, both "b" and "c" are "0", so the graph is of the form y = ax^2 and we already now know that "a" is "2". So, that's all you have to do. The answer is D) 2.

OpenStudy (anonymous):

And for verification: y = 2 = 2(1^2) y = 8 = 2(2^2)

OpenStudy (anonymous):

y=2*x^2

OpenStudy (anonymous):

All good now, @katlin95 ?

OpenStudy (anonymous):

hold on let me find my answer stay here for a sec,

OpenStudy (anonymous):

1/2 is my answer?

OpenStudy (anonymous):

That was one way to get the answer, the graph inspection and here's another, more analytical: y = 8 = a(2^2) = (a)(4) a = 8/4 = 2

OpenStudy (anonymous):

I keep telling you that the answer is "2".

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition

OpenStudy (anonymous):

ohhh lol sowwy

OpenStudy (anonymous):

thanks bud:)

OpenStudy (anonymous):

np! :-)